# Transolver is a Linear Transformer: Revisiting Physics-Attention through the Lens of Linear Attention

### v2

@@ Introduction
We also conduct experiments to investigate the impact of normalization dimension on model performance. Previous works present multiple combinations of normalization dimensions for $\varphi(\bm{Q})$ and $\psi(\bm{K})$ (shen2021efficient; haoGNOTGeneralNeural2023). In this section, compare different combinations of Softmax normalization. Here, we assume $\varphi(\bm{Q}),\psi(\bm{K})\in\mathbb{R}^{N\times M}$. As shown in Table 6, applying Softmax normalization to $\varphi(\bm{Q})$ and $\psi(\bm{K})$ on dimensions $M$ and $N$ performs best, because it maintains the row normalization property of the attention matrix $\varphi(\bm{Q})\psi^{\top}(\bm{K})$.
@@ A. Proof of Theorem 1
In this section, we present the complete proof of Theorem 1, following the proof techniques proposed by DBLP:conf/icml/ChenW24 (DBLP:conf/icml/ChenW24).
@@ B. Dataset and Implementation Details
Airfoil (DBLP:journals/jmlr/LiHLA23) Structure Grid 11,271
Pipe (DBLP:journals/jmlr/LiHLA23) Structure Grid 16,641
Plasticity (DBLP:journals/jmlr/LiHLA23) Structure Grid 3,131
NS (liFourierNeuralOperator2021) Cartesian Grid 4,096
Darcy (liFourierNeuralOperator2021) Cartesian Grid 7,225
Burgers (wangLatentNeuralOperator2024) Cartesian Grid 16,384
Elasticity (DBLP:journals/jmlr/LiHLA23) Unstructure Grid 972
AirfRANS (bonnet2022airfrans) Unstructure Grid 32,000
Shape-Net Car (DBLP:journals/tog/ShapeCar) Unstructure Grid 32,186
@@
Airfoil (DBLP:journals/jmlr/LiHLA23) In this benchmark, the input airfoil geometries are discretized onto a structured grid of size $221\times 51$. The input is the coordinates of each grid point ($221\times 51\times 2$) and the output is the Mach number at each grid point ($221\times 51\times 1$). All geometries are generated by deforming the baseline airfoil NACA-0012. We use 1,000 samples for training and 200 samples for testing.
@@
Pipe (DBLP:journals/jmlr/LiHLA23) In this benchmark, the pipe geometry is discretized onto a structured grid of size $129\times 129$. The input consists of the coordinates of each grid point ($129\times 129\times 2$), and the output is the velocity at each point ($129\times 129\times 1$). The pipe geometries are generated by controlling the centerline of the pipe. A total of 1,000 samples are used for training and 200 samples for testing.
@@
Darcy (liFourierNeuralOperator2021) In this benchmark, the process is discretized into $421\times 421$ Cartesian grid and then downsampled to a resolution of $85\times 85$ for the main experiments. The input consists of the coordinates and the porous medium of each point ($85\times 85\times(2+1)$), while the output is the pressure at each point ($85\times 85\times 1$). For training, 1000 samples are used, 200 samples are generated for testing, and different cases contain different medium structures.
@@
Navier-Stokes (NS) (liFourierNeuralOperator2021) In this benchmark, the flow field is discretized on a $64\times 64$ Cartesian grid. The input consists of the velocity fields from the past 10 frames, with a shape of $64\times 64\times 10\times 1$, and the output is the velocity fields for the next 10 frames with a shape of $64\times 64\times 10\times 1$. The model predicts one frame at once and recursively generates the full 10-frame sequence in an autoregressive manner. For training, 1000 samples with different initial conditions are used, and 200 additional samples are generated for testing.
@@
Plasticity (DBLP:journals/jmlr/LiHLA23) This benchmark aims to model the future deformation behavior of a plastic material when compressed by a die of arbitrary shape. For each sample, the arbitrarily-shaped die is discretized onto a structured grid of size $101\times 31$. The input consists of the coordinates of each grid point and the corresponding external force, resulting in a tensor of shape $101\times 31\times(2+1)$. The output is the deformation of each point along four directions over the next 20 time steps, resulting in a tensor of shape $101\times 31\times 20\times 4$. For training, 900 samples with various die shapes are used, and 80 additional samples are generated for testing.
@@
Elasticity (DBLP:journals/jmlr/LiHLA23) This benchmark focuses on predicting the internal stress distribution within an elastic material, given its structural configuration discretized into 972 points. Each input sample is represented as a tensor of size $972\times 2$, where each row corresponds to the 2D coordinates of a discretized point. The output provides the stress value at each of these locations, formatted as a tensor of shape $972\times 1$. For training, 1000 structurally varied samples are utilized, with an additional 200 samples allocated for testing.
@@
Computer Cost
Transolver 32.38 52.17 9.10 46.16 20.87 0.76
@@
AirfRANS (bonnet2022airfrans) It’s a dataset for studying the two-dimensional incompressible steady-state Reynolds-Averaged Navier-Stokes equations over airfoils at a subsonic regime and for different angles of attack. Unlike the Airfoil dataset (DBLP:journals/jmlr/LiHLA23), this dataset features more complex geometries, a larger number of grid points, richer input variables, and requires predicting more variables. Specifically, it adopts airfoil shapes from the 4 and 5-digit series defined by the National Advisory Committee for Aeronautics (NACA). Each sample consists of approximately 32,000 unstructured grid points covering both the airfoil surface and the surrounding flow region, associated with various flow conditions such as different angles of attack and Reynolds numbers. The input for each point includes the spatial coordinates, the inlet velocity, the Euclidean distance from the point to the airfoil, and the unit outward-pointing surface normal (set to zero for non-surface points), resulting in an input tensor of shape $32000\times(2+2+1+2)$. The output includes the velocity, pressure, and turbulent kinematic viscosity at each point, with a shape of $32000\times(2+1+1)$. The lift coefficient ($C_{L}$) and the Spearman’s correlation coefficient ($\rho_{L}$) can be derived from the predicted outputs. However, following the experimental setup proposed by Transolver (wu2024transolver), we mainly focus on evaluating the pressure field in the surrounding volume (Volume), the surface pressure field on the airfoil (Surface), as well as the lift coefficient ($C_{L}$) and the Spearman’s correlation coefficient ($\rho_{L}$). By varying airfoil geometries, Reynolds numbers, and angles of attack, the AirfRANS dataset provides 1,000 samples in total, with 720 used for training, 80 for validation and 200 for testing.
@@
Shape-Net Car (DBLP:journals/tog/ShapeCar) This dataset simulates the physical fields around and on the surface of 889 different car shapes, each traveling at a speed of 72 km/h. Specifically, the car surface and its surrounding domain are discretized into 32,186 unstructured grid points. The input for each point includes the point position, signed distance function, and normal vector, resulting in an input shape of $32,186\times(3+1+3)$. The output consists of the velocity and pressure at each point, with a shape of $32,186\times(3+1)$. Similar to AirfRANS (bonnet2022airfrans), we compute the drag coefficient ($C_{D}$) and the Spearman’s correlation coefficient ($\rho_{D}$) based on the predicted physical fields. Following Transolver (wu2024transolver), we focus on assessing the velocity field in the surrounding region (Volume), the surface pressure field (Surface), the drag coefficient ($C_{D}$), and the Spearman’s correlation coefficient ($\rho_{D}$). We use 789 samples for training and 100 samples for testing.
@@
Burgers (wangLatentNeuralOperator2024): This dataset is used for the super-resolution experiments presented in Appendix E. Specifically, we adopt the 1D Burgers’ equation as
@@
where $GP$ means Gaussian Process. Then, both time $t\in[0,1]$ and space $x\in[0,1]$ are discretized into a $128\times 128$ grid with periodic boundary conditions. The initial conditions are sampled from a Gaussian process with periodic length $p=1$ and scaling factor $l=1$. The task is to randomly sample a certain number of points $(x^{\prime},u(x^{\prime}))$ from the domain $(x,t)\in[0,1]\times[0.25,0.75]$ as input, and reconstruct the solution $u(x)$ at the domain $(x,t)\in[0,1]\times[0.25,0.75]$. This setup is referred to as the Completer task in Section 4.2 of the LNO (wangLatentNeuralOperator2024). We use 4096 samples for training, 128 samples for validating and 128 samples for testing.
@@
The detailed hyperparameters are listed in Table 8. To ensure a fair comparison, we follow the experimental settings of Transolver (wu2024transolver) for the six standard PDE benchmarks, AirfRANS, and Shape-Net Car. For the Burgers dataset, we adopt the settings from LNO (wangLatentNeuralOperator2024). All models are trained three times on a NVIDIA RTX 4090 GPU. Following Transolver, we adopt a single convolutional layer as a preprocessing step for datasets defined on structured and Cartesian grids, except for the NS dataset, as it corresponds to an unsteady fluid flow problem.
@@ C. Fully Experimental Results
GraphSage (DBLP:conf/nips/HamiltonYL17) 0.0798 0.1254 0.4333 0.9707 0.1192 0.2359 0.2538 0.9894
PointNet (DBLP:conf/cvpr/QiSMG17) 0.0838 0.1403 0.3836 0.9806 0.2021 0.4649 0.4425 0.9784
GraphUNet (DBLP:conf/icml/GaoJ19) 0.0538 0.1168 0.4664 0.9645 0.0979 0.2391 0.3756 0.9816
MeshGraphNet (DBLP:conf/iclr/PfaffFSB21) 0.2789 0.2382 1.7718 0.7631 0.4902 1.1071 0.6525 0.8927
GNO (liNeuralOperatorGraph2020) 0.0833 0.1562 0.4408 0.9878 0.1626 0.2359 0.3038 0.9884
GALERKIN (DBLP:conf/nips/Cao21) 0.0330 0.0972 0.4615 0.9826 0.0577 0.2773 0.3814 0.9821
GNOT (haoGNOTGeneralNeural2023) 0.0305 0.0959 0.3268 0.9865 0.0471 0.3466 0.3497 0.9868
GINO (DBLP:conf/nips/LiKCLKONS0AA23) 0.0839 0.1825 0.4180 0.9645 0.1589 0.2469 0.2583 0.9923
LNO (wangLatentNeuralOperator2024) 0.0825 0.1762 0.7440 0.9399 0.0346 0.0790 0.3732 0.9905
Transolver (wu2024transolver) 0.0122 0.0550 0.1622 0.9904 0.0480 0.2335 0.2438 0.9948
@@ E. Super-Resolution Experiment
Here, $\bm{H}^{(L)}_{K}$ represents the hidden representation obtained after $L$ layers of the LinearNO Block. We selected three baseline models—DeepONet (luLearningNonlinearOperators2021), GNOT (haoGNOTGeneralNeural2023), and LNO (wangLatentNeuralOperator2024)—to compare the reconstructed physical field against the original one using the Relative MAE metric under various sampling rates: 20%, 10%, 5%, 1%, and 0.5%. Since Transolver does not perform the decoupling operation, it cannot be applied to this task. The baseline settings follow those described in Appendix D of LNO (wangLatentNeuralOperator2024). To ensure a fair comparison, our model adopts the same hidden dimension and number of layers as LNO, as shown in Table 8. As shown in the Table 14, our model achieves the best performance for four out of five sampling rates, demonstrating the competitiveness of LinearNO on the super-resolution task.
@@
F. Visualizations

### v3

@@ Introduction
We also conduct experiments to investigate the impact of normalization dimension on model performance. Previous works present multiple combinations of normalization dimensions for $\varphi(\bm{Q})$ and $\psi(\bm{K})$ (Shen et al. 2021; Hao et al. 2023). In this section, compare different combinations of Softmax normalization. Here, we assume $\varphi(\bm{Q}),\psi(\bm{K})\in\mathbb{R}^{N\times M}$. As shown in Table 6, applying Softmax to $\varphi(\bm{Q})$ and $\psi(\bm{K})$ on dimensions $M$ and $N$ performs best, because it maintains the row normalization property of the attention matrix $\varphi(\bm{Q})\psi^{\top}(\bm{K})$.
@@ Conclusion
Acknowledgments
This work was supported by National Key R&D Program 2025YFB3003600 and the Open Fund of National Key Laboratory of Parallel and Distributed Computing (PDL)NO.WDZC20255290101.
@@ A. Proof of Theorem 1
In this section, we present the complete proof of Theorem 1, following the proof techniques proposed by Chen and Wu (2024).
@@ B. Dataset and Implementation Details
Airfoil (2023c) Structure Grid 11,271
Pipe (2023c) Structure Grid 16,641
Plasticity (2023c) Structure Grid 3,131
NS (2021) Cartesian Grid 4,096
Darcy (2021) Cartesian Grid 7,225
Burgers (2024) Cartesian Grid 16,384
Elasticity (2023c) Unstructure Grid 972
AirfRANS (2022) Unstructure Grid 32,000
Shape-Net Car (2018) Unstructure Grid 32,186
@@
Airfoil (Li et al. 2023c) In this benchmark, the input airfoil geometries are discretized onto a structured grid of size $221\times 51$. The input is the coordinates of each grid point ($221\times 51\times 2$) and the output is the Mach number at each grid point ($221\times 51\times 1$). All geometries are generated by deforming the baseline airfoil NACA-0012. We use 1,000 samples for training and 200 samples for testing.
@@
Pipe (Li et al. 2023c) In this benchmark, the pipe geometry is discretized onto a structured grid of size $129\times 129$. The input consists of the coordinates of each grid point ($129\times 129\times 2$), and the output is the velocity at each point ($129\times 129\times 1$). The pipe geometries are generated by controlling the centerline of the pipe. A total of 1,000 samples are used for training and 200 samples for testing.
@@
Darcy (Li et al. 2021) In this benchmark, the process is discretized into $421\times 421$ Cartesian grid and then downsampled to a resolution of $85\times 85$ for the main experiments. The input consists of the coordinates and the porous medium of each point ($85\times 85\times(2+1)$), while the output is the pressure at each point ($85\times 85\times 1$). For training, 1000 samples are used, 200 samples are generated for testing, and different cases contain different medium structures.
@@
Navier-Stokes (NS) (Li et al. 2021) In this benchmark, the flow field is discretized on a $64\times 64$ Cartesian grid. The input consists of the velocity fields from the past 10 frames, with a shape of $64\times 64\times 10\times 1$, and the output is the velocity fields for the next 10 frames with a shape of $64\times 64\times 10\times 1$. The model predicts one frame at once and recursively generates the full 10-frame sequence in an autoregressive manner. For training, 1000 samples with different initial conditions are used, and 200 additional samples are generated for testing.
@@
Plasticity (Li et al. 2023c) This benchmark aims to model the future deformation behavior of a plastic material when compressed by a die of arbitrary shape. For each sample, the arbitrarily-shaped die is discretized onto a structured grid of size $101\times 31$. The input consists of the coordinates of each grid point and the corresponding external force, resulting in a tensor of shape $101\times 31\times(2+1)$. The output is the deformation of each point along four directions over the next 20 time steps, resulting in a tensor of shape $101\times 31\times 20\times 4$. For training, 900 samples with various die shapes are used, and 80 additional samples are generated for testing.
@@
Elasticity (Li et al. 2023c) This benchmark focuses on predicting the internal stress distribution within an elastic material, given its structural configuration discretized into 972 points. Each input sample is represented as a tensor of size $972\times 2$, where each row corresponds to the 2D coordinates of a discretized point. The output provides the stress value at each of these locations, formatted as a tensor of shape $972\times 1$. For training, 1000 structurally varied samples are utilized, with an additional 200 samples allocated for testing.
@@
Computer Cost (GFLOPs) Transolver 32.38 52.17 9.10 46.16 20.87 0.76
@@
AirfRANS (Bonnet et al. 2022) It’s a dataset for studying the two-dimensional incompressible steady-state Reynolds-Averaged Navier-Stokes equations over airfoils at a subsonic regime and for different angles of attack. Unlike the Airfoil dataset (Li et al. 2023c), this dataset features more complex geometries, a larger number of grid points, richer input variables, and requires predicting more variables. Specifically, it adopts airfoil shapes from the 4 and 5-digit series defined by the National Advisory Committee for Aeronautics (NACA). Each sample consists of approximately 32,000 unstructured grid points covering both the airfoil surface and the surrounding flow region, associated with various flow conditions such as different angles of attack and Reynolds numbers. The input for each point includes the spatial coordinates, the inlet velocity, the Euclidean distance from the point to the airfoil, and the unit outward-pointing surface normal (set to zero for non-surface points), resulting in an input tensor of shape $32000\times(2+2+1+2)$. The output includes the velocity, pressure, and turbulent kinematic viscosity at each point, with a shape of $32000\times(2+1+1)$. The lift coefficient ($C_{L}$) and the Spearman’s correlation coefficient ($\rho_{L}$) can be derived from the predicted outputs. However, following the experimental setup proposed by Transolver (Wu et al. 2024), we mainly focus on evaluating the pressure field in the surrounding volume (Volume), the surface pressure field on the airfoil (Surface), as well as the lift coefficient ($C_{L}$) and the Spearman’s correlation coefficient ($\rho_{L}$). By varying airfoil geometries, Reynolds numbers, and angles of attack, the AirfRANS dataset provides 1,000 samples in total, with 720 used for training, 80 for validation and 200 for testing.
@@
Shape-Net Car (Umetani and Bickel 2018) This dataset simulates the physical fields around and on the surface of 889 different car shapes, each traveling at a speed of 72 km/h. Specifically, the car surface and its surrounding domain are discretized into 32,186 unstructured grid points. The input for each point includes the point position, signed distance function, and normal vector, resulting in an input shape of $32,186\times(3+1+3)$. The output consists of the velocity and pressure at each point, with a shape of $32,186\times(3+1)$. Similar to AirfRANS (Bonnet et al. 2022), we compute the drag coefficient ($C_{D}$) and the Spearman’s correlation coefficient ($\rho_{D}$) based on the predicted physical fields. Following Transolver (Wu et al. 2024), we focus on assessing the velocity field in the surrounding region (Volume), the surface pressure field (Surface), the drag coefficient ($C_{D}$), and the Spearman’s correlation coefficient ($\rho_{D}$). We use 789 samples for training and 100 samples for testing.
@@
Burgers (Wang and Wang 2024): This dataset is used for the super-resolution experiments presented in Appendix E. Specifically, we adopt the 1D Burgers’ equation as
@@
where $GP$ means Gaussian Process. Then, both time $t\in[0,1]$ and space $x\in[0,1]$ are discretized into a $128\times 128$ grid with periodic boundary conditions. The initial conditions are sampled from a Gaussian process with periodic length $p=1$ and scaling factor $l=1$. The task is to randomly sample a certain number of points $(x^{\prime},u(x^{\prime}))$ from the domain $(x,t)\in[0,1]\times[0.25,0.75]$ as input, and reconstruct the solution $u(x)$ at the domain $(x,t)\in[0,1]\times[0.25,0.75]$. This setup is referred to as the Completer task in Section 4.2 of the LNO (Wang and Wang 2024). We use 4096 samples for training, 128 samples for validating and 128 samples for testing.
@@
The detailed hyperparameters are listed in Table 8. To ensure a fair comparison, we follow the experimental settings of Transolver (Wu et al. 2024) for the six standard PDE benchmarks, AirfRANS, and Shape-Net Car. For the Burgers dataset, we adopt the settings from LNO (Wang and Wang 2024). All models are trained three times on a NVIDIA RTX 4090 GPU. Following Transolver, we adopt a single convolutional layer as a preprocessing step for datasets defined on structured and Cartesian grids, except for the NS dataset, as it corresponds to an unsteady fluid flow problem.
@@ C. Fully Experimental Results
GraphSage (2017) 0.0798 0.1254 0.4333 0.9707 0.1192 0.2359 0.2538 0.9894
PointNet (2017) 0.0838 0.1403 0.3836 0.9806 0.2021 0.4649 0.4425 0.9784
GraphUNet (2019) 0.0538 0.1168 0.4664 0.9645 0.0979 0.2391 0.3756 0.9816
MeshGraphNet (2020) 0.2789 0.2382 1.7718 0.7631 0.4902 1.1071 0.6525 0.8927
GNO (2020) 0.0833 0.1562 0.4408 0.9878 0.1626 0.2359 0.3038 0.9884
GALERKIN (2021) 0.0330 0.0972 0.4615 0.9826 0.0577 0.2773 0.3814 0.9821
GNOT (2023) 0.0305 0.0959 0.3268 0.9865 0.0471 0.3466 0.3497 0.9868
GINO (2023d) 0.0839 0.1825 0.4180 0.9645 0.1589 0.2469 0.2583 0.9923
LNO (2024) 0.0825 0.1762 0.7440 0.9399 0.0346 0.0790 0.3732 0.9905
Transolver (2024) 0.0122 0.0550 0.1622 0.9904 0.0480 0.2335 0.2438 0.9948
@@ E. Super-Resolution Experiment
Here, $\bm{H}^{(L)}_{K}$ represents the hidden representation obtained after $L$ layers of the LinearNO Block. We selected three baseline models—DeepONet (Lu et al. 2021), GNOT (Hao et al. 2023), and LNO (Wang and Wang 2024)—to compare the reconstructed physical field against the original one using the Relative MAE metric under various sampling rates: 20%, 10%, 5%, 1%, and 0.5%. Since Transolver does not perform the decoupling operation, it cannot be applied to this task. The baseline settings follow those described in Appendix D of LNO (Wang and Wang 2024). To ensure a fair comparison, our model adopts the same hidden dimension and number of layers as LNO, as shown in Table 8. As shown in the Table 14, our model achieves the best performance for four out of five sampling rates, demonstrating the competitiveness of LinearNO on the super-resolution task.
@@
F. Equivalence of Cross-Attention with Learnable Tokens to Linear Attention
Cross-attention is originally introduced to model interactions between two sequences, allowing one sequence to attend to relevant information in the other. However, in cross-attention with learnable tokens, most existing works treat these tokens as coordinate anchors in a latent space, which are used to compress the original sequence (Wang and Wang 2024; Serrano et al. 2024; Alkin et al. 2024). Taking the Physics-Cross-Attention module in LNO (Wang and Wang 2024) as an example, the operation can be formulated as follows:
$\bm{S}_{M}$ $=\text{Physics-Cross-Attention}(\bm{Z}_{M},\bm{H}_{N},\bm{\hat{H}}_{N})$ (31)
$=\text{Softmax}\left(\frac{1}{\sqrt{d_{h}}}\bm{Z_{M}}{{W}_{q}}{{W}_{k}}^{\top}\bm{H}_{N}^{\top}\right)\bm{\hat{H}}_{N}{W}_{v}$
$=\text{Softmax}\left({W}_{1}\bm{H}_{N}^{\top}\right)\bm{\hat{H}}_{N}{W}_{v}$
$=\text{Softmax}((\text{Linear}^{(K)}(\bm{H}_{N})^{\top})\text{Linear}^{(V)}(\bm{\hat{H}}_{N})$
$\bm{S}_{M}^{\prime}=\text{Self-Attention}\circ\dots\circ\text{Self-Attention}(\bm{S}_{M})$ (32)
$\bm{H}_{N}^{\prime}$ $=\text{Physics-Cross-Attention}(\bm{H}_{N},\bm{Z}_{M},\bm{S}^{\prime}_{M})$ (33)
$=\text{Softmax}\left(\frac{1}{\sqrt{d_{h}}}\bm{H}_{N}{W}_{q}^{\prime}{W}_{k}^{\top}\bm{Z}_{M}^{\top}\right)\bm{S}^{\prime}_{M}{W}_{v}^{\prime}$
$=\text{Softmax}\left(\bm{H}_{N}{W}_{2}^{\top}\right)\bm{S}^{\prime}_{M}{W}_{v}^{\prime}$
$=\text{Softmax}(\text{Linear}^{(Q)}(\bm{H}_{N}))\bm{S}^{\prime}_{M}{W}_{v}^{\prime}$
where $\bm{Z}_{M}=\{\bm{z}_{i}\}_{i=1}^{M}\in\mathbb{R}^{M\times d_{h}}$ denotes the learnable tokens. $W_{q},W_{k},W_{v},W^{\prime}_{q},W^{\prime}_{k},$ and $W^{\prime}_{v}$ are learnable matrices in $\mathbb{R}^{d_{h}\times d_{h}}$. By the associativity of matrix multiplication, the matrices $\frac{1}{\sqrt{d_{h}}}{Z_{M}}{W}_{q}{W}_{k}^{\top}$ and $\frac{1}{\sqrt{d_{h}}}{W}_{q}^{\prime}{W}_{k}^{\top}{Z_{M}}^{\top}$ can be merged into two learnable matrices, denoted as $W_{1}$ and $W_{2}$. In LNO, $\bm{H}_{N}$ contains only coordinate information, while $\bm{\hat{H}}_{N}$ includes additional information beyond coordinates, such as the porous medium of each point in the Darcy dataset.
If the compression is based solely on spatial coordinates, $\bm{H}_{N}$ only contains coordinate information; otherwise, when full point-wise information is used, $\bm{H}_{N}=\bm{\hat{H}}_{N}$. Whether $\bm{H}_{N}$ is equal to $\bm{\hat{H}}_{N}$ is an implementation detail and does not materially affect our discussion. Therefore, we can construct equivalent feature mappings $\varphi(\bm{Q})$, $\psi(\bm{K})$, $\bm{V}$ and $\mathcal{G}$ as in Eq.(34),
$\varphi(\bm{Q})=\text{Softmax}(\text{Linear}^{(Q)}(\bm{H}_{N}))$ (34)
$\psi^{\top}(\bm{K})=\text{Softmax}(\text{Linear}^{(K)}(\bm{H}_{N})^{\top})$
$\bm{V}=\text{Linear}^{(V)}(\bm{\hat{H}}_{N})$
$\mathcal{G}=W^{\prime}_{v}\circ\text{Self-Attention}\circ\dots\circ\text{Self-Attention}$
Overall, cross-attention with learnable tokens can also be regarded as a special case of linear attention.
G. Visualizations